7. Suppose that you have been hired as an economist by OPEC and given the following schedule showing the world demand for oil:Your advice is needed on the following Price(S/barrel) Quantity demanded(millions of barrels/day) 10 20 30 40 50 60 50 40 30 20 questions: a) If the price rises from $20 to $30 a barrel, will the total revenue from oil sales increase or decrease? b) What will happen to total revenue if the supply of oil is decreased further and the price rises from $30 to S40 a barrel? c) What is the price that will achieve the highest total revenue? (the price may be not shown in the table) d) What quantity of oil will be sold at the answer to (c)? e) What are the values of the price elasticity of demand for price changes of S10 a barrel at average prices of $15 and $45? f) What is the elasticity of demand that maximizes total revenue? g) Over what range of prices is the demand for oil inelastic?
media%2Ff6a%2Ff6a277f4-9134-42ff-97ee-ca
media%2F991%2F991ba2c0-878f-42a2-9c20-6d
media%2Fc87%2Fc8717dd1-b9aa-4046-9455-40
Answer #1

Answer 7 - We have been given a d table which shows price and quantity demanded of oil. We know that that total revenue is equal to price times quantity.

TR = P * Q

Price ($/barrel) Quantity Demanded (Millions of barrels per day) Total revenue (TR=P*Q) Marginal Revenue (Change in TR/change in quantity sold)
10 60 600
20 50 1000 400/10 = 40
30 40 1200 200/10 = 20
40 30 1200 0/10 = 0
50 20 1000 -200/10 = -20

(a) Look at the table, If price rise from $20 per barrel to $30 per barrel then total revenue will rise. Total revenue is $1000 million if price is $20 per barrel and TR is $1200 million if price is $30 per barrel.

(b) Total revenue will be constant at $1200 million if price rise from $30 to $40 per barrel. In case of both price total revenue is equal to $1200 millions. Thus TR will not change.

(c) OPEC will have highest revenue of $1200 millions. OPEC will have maximum total revenue when marginal revenue reaches to ''zero''. Revenue maximization condition is MR = 0. MR is equal to zero at price $40 per barrel.

(d) At price of $40 OPEC will sell 30 million barrel of oil in the world market. OPEC will earn maximum revenue at this level of quantity demanded.